A review of the ledger of Weakly Service Co. at December 31, 2017, produces the following data pertaining to the preparation of annual adjusting entries: (a) Notes Payable $80,000: This is a 9-month note, dated September 1, 2017, with a 9% interest rate. (b) Prepaid Rent $648,000. The company rents offices throughout the Midwest. During 2017 it signed 10 leases as shown below: Date Term (in months) Monthly Rent Number of Leases Sept. 1 8 $4,500 4 Nov. 1 12 7,000 6 (c) Unearned Service Revenue $171,000. During 2017 the company entered into 13 monthly service contracts with clients. The clients prepaid for the services to be provided over the contract period in an even manner. Date Service period (in months) Amount Per Contract Number of Contracts Aug. 1 9 $12,600 8 Oct. 1 6 15,000 5 Prepare the adjusting entries at December 31, 2017. (Credit account titles are automatically indented when the amount is entered. Do not indent manually.)
Answer #1

Adjusting entries are provided below and computations are shown below each entry in brackets:

Date Particulars Dr Cr
12/31/17 Interest expense 2,400
Interest payable 2,400
($80,000*0.09*4/12 = 2400)
12/31/17 Rent expense 156,000
Prepaid rent 156,000
(4*4500*4 = 72000)
(2*7000*6 = 84000)
(72000+84000 = 156000)
12/31/17 Unearned service revenue 93,500
Service revenue 93,500
(5/9*12600*8 = 56000)
(3/6*15000*5 = 37500)
(56000+37500 = 93500)